OpenStudy (anonymous):

Probability: A box of crayons contains 3 shades of red, 5 shades of blue, and 2 shades of green. if a child chooses a crayon at random, find the probablitiy of choosing the following gree crayon red crayon blue crayon crayon tht is not red red or blue crayon red or green crayon

OpenStudy (anonymous):

The probability of choosing a green crayon is (number of green crayons)/(total number of crayons), or 2/(2+3+5) = 2/10 = 1/5. Can you use that same method to solve for red and blue crayons?

OpenStudy (anonymous):

dont have time

OpenStudy (anonymous):

? I'm not here to do your homework for you. If you don't have time to do your homework, you should have set aside more time to work on it.

OpenStudy (anonymous):

really i dont

OpenStudy (anonymous):

Well, next time I'm sure you'll set aside more time for your homework then.

OpenStudy (amistre64):

the time taken to post it on here could have finished it up .... right?

OpenStudy (anonymous):

nm

OpenStudy (amistre64):

http://openstudy.com/code-of-conduct just so you know, as long as you dont expect us to hand out answers to you; its because we ar enot that kind of website. if you have time to learn it, we are here to help you out if you need it thank